# Codimension-Two Spiral Spin-Liquid in the Effective Honeycomb-Lattice Compound Cs3Fe2Cl9

### v1

@@ III Magnetization measurements
Figure S1: Refinement result of the powder neutron diffraction data measured on POWGEN at $T$ = 15 K. Data points are shown as red circles. The calculated pattern is shown as the black solid line. The vertical bars indicate the positions of the structural Bragg peaks for Cs3Fe2Cl9. The blue line at the bottom shows the difference of measured and calculated intensities.
@@ IV Neutron diffraction experiments on POWGEN
Figure S1 summarizes the refinement result of the neutron diffraction pattern collected at $T=15$ K. No secondary reflections are observed in the diffraction pattern, which confirms the phase purity of our sample. Refined crystal structure parameters are listed in Table S1. Due to the lacking of neutron diffraction data in the high-$Q$ region, we cannot reliable fit the thermal parameters. Therefore, a uniform thermal factor of $B_{\textrm{iso}}=0.1$2 is assumed for all atoms. Figure S2 plots the temperature evolution of the diffraction pattern. Below $T_{\textrm{N}}\sim 5.5$ K, magnetic Bragg peaks belonging to $\bm{q}=(\frac{1}{2},0,0)$ are observed. This transition temperature is consistent with that observed in magnetic susceptibility.
@@
Table S1: Refined crystal structure parameters for Cs3Fe2Cl9 at 15 K. The space group is $P6_{3}/mmc$ with lattice constants $a=7.1784(1)$ and $c=17.6374(3)$. A uniform thermal parameter of $B_{\textrm{iso}}=0.1$2 is assumed for all atoms. The goodness-of-fit parameters are $R_{p}=18.6\%$ and $R_{wp}=10.2\%$.
@@
Table S2: Refined crystal structure parameters for Cs3Fe2Cl9 at 1.6 K. The space group is $P6_{3}/mmc$ with lattice constants $a=7.1784(1)$ and $c=17.6369(2)$. A uniform thermal parameter of $B_{\textrm{iso}}=0.1$2 is assumed for all atoms. The goodness-of-fit parameters are $R_{p}=15.2\%$ and $R_{wp}=11.3\%$.
@@ V Inelastic neutron scattering experiments on CNCS
Inelastic neutron scattering (INS) experiments on powder sample of Cs3Fe2Cl9 were performed on CNCS Ehlers et al. (2011) at the SNS of the ORNL. About 5 g powder was sealed in an aluminum can in a helium filled glovebox. An orange cryostat was utilized to reach a base temperature of 5 K. Measurements were taken with incident neutron energies of $E_{i}=6.59$, 3.32, and 1.0 meV in the high flux chopper configuration at $T=2$ and 10 K. For each measuring condition, data were collected on an empty can and subtracted as background. Data reduction and projection were performed using the Mslice program in DAVE Azuah et al. (2009).
Figure S3: INS spectra for Cs3Fe2Cl9 measured on CNCS with an incident neutron energy of $E_{i}=1.0$ meV at $T=2$ K. Data were collected for approximately 8 hrs with the source operating at a power of 1.4 MW.
Figure S3 presents the INS spectra for Cs3Fe2Cl9 measured with an incident neutron nergy of $E_{i}=1.0$ meV at $T=2$ K. An excitation gap of $\sim 0.2$ meV is resolved, which is consistent with the existence of uniaxial SIA in Cs3Fe2Cl9.
@@ VI Diffuse neutron scattering experiments on CORELLI
Figure S4 compares the diffuse scattering patterns calculated using the SCGA method for the fitted $J_{123}$-$D_{z}$ model at $T=5$ and 6 K. As explained in the main text, the SCGA method is a mean-field-based theory and may underestimate the critical scattering in systems with relatively strong SIA. If we assume $T=6$ K as that in experiments, the calculated scattering pattern will exhibit less contrast at momentum transfers over and out of the spiral surface. Therefore, a reduced temperature of $T=5$ K is assumed in Fig. 3 of the main text to compensate for the underestimated critical scattering. On the other hand, the comparison between the calculations at $T=5$ and 6 K also reveals that the scattering intensity between the integer-$l$ planes is reduced at lower temperatures. This temperature dependence confirms thermal excitations as the origin of the interplanar scattering.
Figure S4: Comparison between the diffuse scattering patterns calculated using the SCGA method for the fitted $J_{123}$-$D_{z}$ model at $T=5$ K (a-f) and 6 K (g-l). Panels (a-f) is a reproduction of Fig. 3 in the main text.
Figure S5: (a) Calculated diffuse scattering pattern in the ($h,~{}k,~{}0$) plane using the fitted $J_{123}$-$D_{z}$ model at $T=6$ K. (b) Temperature evolution of the calculated spin correlation function $\langle\bm{S}\cdot\bm{S}\rangle$ along ($h$, $h$, 0) around (0.5, 0.5, 0). The integration area is outlined in panel (a) by a dashed rectangle. Black dots indicate the peak positions at the corresponding temperatures through Gaussian fits. (c) Temperature evolution of the calculated spin correlations $\langle S_{x}\cdot S_{x}\rangle$ (red triangles), $\langle S_{z}\cdot S_{z}\rangle$ (yellow circles), and their ratio $\langle S_{x}\cdot S_{x}\rangle/\langle S_{z}\cdot S_{z}\rangle$ (black dots). The integration area is outlined in panel (a) by a dashed-line hexagon. Red line over the $\langle S_{x}\cdot S_{x}\rangle/\langle S_{z}\cdot S_{z}\rangle$ data is the fitting curve to the Arrhenius law $A_{1}\exp(-\frac{T-T_{N}}{\Delta_{\textrm{SIA}}})+A_{2}$ with fitted parameters of $A_{1}=-0.52(1)$, $\Delta_{\textrm{SIA}}=4.03$ K, and $A_{2}=0.90(1)$.
VII Effects of the uniaxial anisotropy on the SSL
Classical Monte Carlo simulations for the fitted $J_{123}$-$D_{z}$ model were performed to study the effects of the SIA on the SSL in Cs3Fe2Cl9. As a reference, the diffuse scattering pattern at $T=6$ K calculated by the SCGA method is shown in Fig. S5. Through classical Monte Carlo simulations that incorporate critical scattering near $T_{N}$, Fig. S5(b) summarizes the temperature evolution of the calculated scattering intensity along ($h+\frac{1}{2}$, $h+\frac{1}{2}$, 0), where the integration range is outlined by dashed rectangle in Fig. S5(b). For the fitted $J_{123}$-$D_{z}$ model, $T_{N}$ is found to be $\sim 5.5$ K. The separated peaks at $T/T_{N}>1$ correspond to thermally stabilized spiral correlations in the SSL state, while their gradually reduced separation with decreasing $T$ reveals the second-order character of the transition into the LRO. Figure S5(d) compares the in-plane ($x$) and out-of-plane ($z$) components of the calculated spin correlations as a function of $T$. Due to the existence of uniaxial SIA, the evolution of $\langle S_{x}\cdot S_{x}\rangle$ is nonmonotonic, and the ratio between the in-plane and out-of-plane components exhibits an Arrhenius-like behavior in a wide temperature regime.
VIII Neutron diffraction experiments on WAND2
@@
Single crystal neutron diffraction measurements were performed using the WAND2 diffractometer Frontzek et al. (2018) at the High Flux Isotope Reactor HFIR with $\lambda=1.486$. A crystal with mass of $\sim 7.5$ mg was aligned with the $c$ axis vertical, and then sealed in an aluminum can with helium exchange gas. A vertical field cryomagnet was used, providing a base temperature of 1.5 K and a maximum field of $H=6$ T. Measurements were performed in phase I at $T$ = 1.5 K and no applied field rotating through 180° in 0.1° steps over 20 h. Measurements in phase II were made at $T$ = 1.5 K and $H=3.2$ T by rotating through 180° in 0.1° steps over 20 h. In phase III at $T=1.5$ K at $H=4.5$ T the sample was rotated through 90° in 0.1° steps over 10 h. In phase IV, at $T=1.5$ K and $H=5.5$ T the sample was rotated through 90° in 0.1° steps over 10 h. In phase V, at $T=4$ K and $H=5.5$ T the sample was rotated through 90° in 0.1° steps over 10 h. In phase VI, at $T=5.05$ K and $H=5.5$ T the sample was rotated over 90° in 0.1° steps with a total measurement time of 26 h.
@@
Figure S6: Single crystal neutron diffraction pattern for Cs3Fe2Cl9 in phases (a) III, (b) IV, and (c) V in the ($h$,$k$,0) plane. The experimental conditions are listed in the text. The dashed rectangle emphasizes the area in which the relative intensity of the $\bm{q}_{\textrm{III}}$ and $2\bm{q}_{\textrm{III}}$ reflections are compared in the text.
@@
Figure S6 compares the neutron diffraction patterns in phases III, IV, and V. As explained in the main text, similar diffraction patterns were observed in these three phases. Especially, their magnetic propagation vector stays at the same $\bm{q}_{\textrm{III}}=\frac{5}{6}\times(\frac{1}{3},\frac{1}{3},0)$, with weak $2\bm{q}_{\textrm{III}}$ higher harmonics observed along the Brillouin zone boundary. Despite their similarities, careful comparisons of the diffraction patterns reveal some differences among these three phases. As emphasized by the dashed rectangle in each panel, the intensity of the $2\bm{q}_{\textrm{III}}$ reflection in phase III is weaker than that in phase IV, which suggests higher field-induced magnetization in phase IV. Comparing phases IV and V, the $\bm{q}_{\textrm{III}}$ reflection becomes weaker in phase V while the $2\bm{q}_{\textrm{III}}$ reflection stays almost unchanged. This evolution is consistent with a transition from a spiral structure in phase IV to a spin density wave (SDW) structure in phase V.
IX Refinements of the nuclear and magnetic structures
@@
Figure S8 summarizes the fitting results of the magnetic reflections collected on WAND2 in phases I (a), III (b), VI (c), and V (d). The measuring conditions are listed in the main text. In phase VI, only two nonequivalent magnetic reflections were obtained for each magnetic domain, which does not allow a reliable analysis of the magnetic structure. For phase I, the collinear magnetic structure as shown in Fig. 1(a) of the main text was assumed in the fits. The fitted magnitude of the ordered magnetic moment is 2.9(2) $\mu\rm{{}_{B}}$, with goodness-of-fit parameters $R_{F2}=20.3\%$ and $R_{F2w}=24.4\%$. For phases III-V, a helical magnetic structure was assumed based on our classical Monte Carlo simulations. The magnitudes of the in-plane magnetic moment, $M_{\parallel}$, and out-of-plane magnetic moment, $M_{\perp}$, were treated as fitting parameters. For phase III, the fitted magnitudes are $M_{\parallel}=1.9\pm 1.5$ and $M_{\perp}=1.9\pm 0.4$ $\mu\rm{{}_{B}}$, with goodness-of-fit parameters $R_{F2}=13.0\%$ and $R_{F2w}=14.7\%$. For phase IV, the fitted magnitudes are $M_{\parallel}=2.1\pm 1.6$ and $M_{\perp}=1.8\pm 0.5$ $\mu\rm{{}_{B}}$, with $R_{F2}=9.0\%$ and $R_{F2w}=10.1\%$. For phase V, the fitted magnitudes are $M_{\parallel}=1.8\pm 4.1$ and $M_{\perp}=1.2\pm 0.8$ $\mu\rm{{}_{B}}$, with $R_{F2}=3.2\%$ and $R_{F2w}=3.7\%$. Considering the standard deviations, the fitted magnitude of the in-plane moment $M_{\parallel}$ is consistent with our classical Monte Carlo simulations, where phases III and IV exhibit an elongated helical structure, and phase V exhibits a spin density wave structure with only the $S_{z}$ components.
X Classical Monte Carlo simulations
Classical Monte Carlo simulations for the $J_{123}$-$D_{z}$ model
$\mathcal{H}=\sum_{\langle ij\rangle\in n}J_{n}\bm{S}_{i}\cdot\bm{S}_{j}+D_{z}(S_{z})^{2}$ (S1)
@@ X Classical Monte Carlo simulations
and its perturbed derivatives were performed using the SpinMC code that implements the single spin flip Metropolis algorithm bue . Unless alternately specified, a $12\times 12\times 4$ supercell with 2304 spins was employed in simulations. Observables including the heat capacity and magnetization were averaged over $10^{6}$ of measurement sweeps after $3\times 10^{5}$ thermalization sweeps, where each sweep represents 2304 attempted spin flips at randomly selected sites. Magnetic structure factors were calculated through fast Fourier transform using the FFTW package fft . The parallel tempering algorithm was utilized to facilitate thermal equilibrium, which was performed simultaneously over 128 replicas on 128 cores with a geometric series of temperatures between 1 and 10 K. After every 10 Monte Carlo sweeps, a replica exchange was attempted and the successful exchange rates are above $\sim 25\%$ for all the neighboring replicas.
X.1 The $J_{123}$ model
Figure S9 summarizes the simulation results for the $J_{123}$ model with no SIA. The coupling strengths are $J_{1}=-0.238$, $J_{2}=0.085$, $J_{3}=0.053$ meV as fitted from the INS experimental data. In zero field, the $J_{123}$ model exhibits a long-range order transition at $T_{N}\sim 3.5$ K, which is lower than that observed in experiments. The phase diagram shown in Fig. S9(a) mainly consists of two phases. Magnetic structure factors shown in Figs. S9(b-d) with $\bm{q}\sim\frac{4}{3}\times(\frac{1}{3},\frac{1}{3},0)$ in low magnetic fields and $\bm{q}=(\frac{1}{2},\frac{1}{2},0)$ in high magnetic fields, which is very different from the experimental observations.
Figure S9: (a) Theoretical phase diagram for the $J_{123}$ model calculated by the classical Monte Carlo simulations. Pseudocolor corresponds to the value of heat capacity in arbitrary units (arb. units). (b-d) Magnetic structure factors $\langle\bm{S}\cdot\bm{S}\rangle$ calculated at $T=1$ K in a field of (b) 0 T, (c) 5 T, and (d) 10 T. Calculations of the magnetic structure factor were performed on a $18\times 18\times 4$ supercell with $3\times 10^{5}$ thermalization sweeps and $5\times 10^{5}$ measurement sweeps.
Figure S10: (a) Theoretical phase diagram for the $J_{123}+D_{z}$ model calculated by the classical Monte Carlo simulations. An easy-axis SIA of $D_{z}=-0.037$ meV is added to the $J_{123}$ model considered in Fig. S9. Pseudocolor corresponds to the value of heat capacity in arbitrary units (arb. units). (b-d) Magnetic structure factors $\langle\bm{S}\cdot\bm{S}\rangle$ calculated at $T=1$ K in a field of (b) 0 T, (c) 6 T, and (d) 10 T.
Figure S11: Exchange paths for the perturbational couplings $J_{4}$ $J_{10}$ that are considered in our Monte Carlo simulations.
X.2 The $J_{123}+D_{z}$ model
Figure S10 summarizes the simulation results for the $J_{123}+{D_{z}}$ model that considers isotropic exchange interactions up to the third neighbors and a uniaxial SIA. The coupling strengths are $J_{1}=-0.238$, $J_{2}=0.085$, $J_{3}=0.053$, and $D_{z}=-0.037$ meV as fitted from the INS experimental data. In zero field, a magnetic long-range order transition is observed at $T_{N}\sim 5.8$ K, which is close to the experimental observation. Although the phase diagram shown in Fig. S10(a) mainly consists of two phases as that in Fig. S9(a), magnetic structure factors shown in Figs. S10(b-d) reveal that the propagation vector $\bm{q}$ stays at $(\frac{1}{2},\frac{1}{2},0)$ r.l.u. in both phases. This calculation indicates that the experimentally observed $\bm{q}_{\textrm{I}}=(\frac{1}{2},\frac{1}{2},0)$ r.l.u. in phase I is due to the existence of uniaxial SIA, yet the stabilization of the field-induced phases III, IV, V, and VI requires further perturbations in the spin Hamiltonian.
Figure S12: (a) Theoretical phase diagram for the $J_{123}+D_{z}+J_{6}$ model calculated by the classical Monte Carlo simulations. A perturbative sixth-neighbor interaction of $J_{6}=0.01$ meV is added to the $J_{123}+D_{z}$ model considered in Fig. S10. Pseudocolor corresponds to the value of heat capacity in arbitrary units (arb. units). (b-d) Magnetic structure factors $\langle\bm{S}\cdot\bm{S}\rangle$ calculated at $T=1$ K in a field of (b) 0 T, (c) 5 T, and (d) 10 T.
X.3 An overview of the perturbation effects
As the theoretical phase diagram for the $J_{123}+D_{z}$ model cannot reproduce the experimental results, extensive classical Monte Carlo simulations were performed to explore possible perturbative terms that may stabilize the experimentally observed additional phases in intermediate fields. By confining the perturbations to a single-term Heisenberg exchange interaction up to the tenth neighbors, we calculated the phase diagram and magnetic structure factors for each perturbed model. The exchange paths for $J_{4}$-$J_{10}$ are indicated in Fig. S11.
Figure S13: (a) Theoretical phase diagram for the $J_{123}+D_{z}+J_{5}$ model calculated by the classical Monte Carlo simulations. A perturbative fifth-neighbor interaction of $J_{5}=-0.008$ meV is added to the $J_{123}-D_{z}$ model considered in Fig. S10. Pseudocolor corresponds to the value of heat capacity in arbitrary units (arb. units). (b-d) Magnetic structure factors $\langle\bm{S}\cdot\bm{S}\rangle$ calculated at $T=1$ K in a field of (b) 0 T, (c) 5 T, and (d) 10 T. Calculations of the magnetic structure factor were performed on a $18\times 18\times 4$ supercell with $3\times 10^{5}$ thermalization sweeps and $5\times 10^{5}$ measurement sweeps. White arrows indicate the positions of weak magnetic Bragg peaks.
Figure S14: Theoretical $M/H$ curves calculated by the classical Monte Carlo simulations for the $J_{123}+D_{z}+J_{5}$ model. The strength of the applied field for each curve is indicated on the left. Data are successively shifted by 0.003 arb. units along the $y$ axis for clarity. The dashed line at $T=1.8$ K marks the lower limit of the measuring temperature in Ref. Ishii et al. (2021).
Through our calculations, it was found that all of the considered perturbations with a strength lower than $\sim 0.01$ meV are able to introduce additional phases in the intermediate field regime, yet their effects on the magnetic structure factors categorize them into two groups. The first group contains antiferromagnetic (AFM) $J_{6}$ and $J_{7}$, or ferromagnetic (FM) $J_{8}$, $J_{9}$, and $J_{10}$ perturbations. However, in zero field, these types of perturbations do not produce the experimentally observed $\bm{q}_{\textrm{I}}=(\frac{1}{2},\frac{1}{2},0)$ in phase I and thus contradicts the experimental observations. The second group contains AFM $J_{4}$ or FM $J_{5}$ perturbations. This group of perturbations stabilize $\bm{q}_{\textrm{I}}=(\frac{1}{2},\frac{1}{2},0)$ in zero field while introducing additional phases similar to that observed in experiments. For both groups of perturbations, a sign reversal eliminates the field-induced phases, leading to phase diagrams similar to that of the unperturbed model shown in Fig. S10(a).
Figure S15: Theoretical heat capacity $C_{p}$ calculated by the classical Monte Carlo simulations for the $J_{123}+D_{z}+J_{5}$ model in the field regime of (a) [0, 3], (b) [4, 7], and (c) [8, 14] T. Data are normalized by the maximal $C_{p}$ value, $C_{p}^{\textrm{max}}$. The strength of the applied field for each curve is indicated on the left. In each panel, data are successively shifted by 1 unit along the $y$ axis for clarity. The dashed line at $T=2.0$ K marks the lower limit of the measuring temperature in Ref. Ishii et al. (2021).
X.4 The $J_{123}+D_{z}+J_{6}$ model
As an illustration for the first group of perturbations, Fig. S10 summarizes the simulation results for the $J_{123}+{D_{z}}+J_{6}$ model that considers a sixth-neighbor AFM perturbation of $J_{6}=0.01$ meV. Comparison of the phase diagrams shown in Fig. S10(a) and Fig. S12(a) reveals that the perturbation of $J_{6}$ causes additional phases in the intermediate field regime between $\sim 4.5$ and 6.5 T. However, as shown in Fig. S12(b-d), the propagation vector of the long-range ordered phase in zero field deviates from $\bm{q}_{\textrm{I}}=(\frac{1}{2},\frac{1}{2},0)$ r.l.u. and thus fails to reproduce the experimental results. Similar effects are observed for the remaining perturbations in the first group, thus excluding them as the main perturbation to the $J_{123}+{D_{z}}$ Hamiltonian.
X.5 The $J_{123}+D_{z}+J_{5}$ model
As discussed in the previous section, AFM $J_{4}$ and FM $J_{5}$ belong to the second group of perturbations and have similar effects on the phase diagram. However, there is one slight difference for these two perturbations: In zero field, AFM $J_{4}$ and FM $J_{5}$ raises and lowers $T_{N}$, respectively. This observation favors FM $J_{5}$ since the experimental $T_{N}\sim 5.4$ K is lower than the theoretically predicted $T_{N}\sim 5.8$ K for the unperturbed $J_{123}+D_{z}$ model.

### v2

@@ III Magnetization measurements
Figure S1: Refinement result of the powder neutron diffraction data measured on POWGEN at $T$ = 15 K. Data points are shown as red circles. The calculated pattern is shown as the black solid line. The vertical bars indicate the positions of the structural Bragg peaks for Cs3Fe2Cl9. The blue line at the bottom shows the difference between measured and calculated intensities.
@@ IV Neutron diffraction experiments on POWGEN
Figure S1 summarizes the refinement result of the neutron diffraction pattern collected at $T=15$ K. No secondary reflections are observed in the diffraction pattern, which confirms the phase purity of our sample. Refined crystal structure parameters are listed in Table S1. Due to the lacking of neutron diffraction data in the high-$Q$ region, we cannot reliably fit the thermal parameters. Therefore, a uniform thermal factor of $B_{\textrm{iso}}=0.1$ Å2 is assumed for all atoms. Figure S2 plots the temperature evolution of the diffraction pattern. Below $T_{\textrm{N}}\sim 5.5$ K, magnetic Bragg peaks belonging to $\bm{q}=(\frac{1}{2},0,0)$ are observed. This transition temperature is consistent with that observed in magnetic susceptibility.
@@
Table S1: Refined crystal structure parameters for Cs3Fe2Cl9 at 15 K. The space group is $P6_{3}/mmc$ with lattice constants $a=7.1784(1)$ and $c=17.6374(3)$ Å. A uniform thermal parameter of $B_{\textrm{iso}}=0.1$ Å2 is assumed for all atoms. The goodness-of-fit parameters are $R_{p}=18.6\%$ and $R_{wp}=10.2\%$.
@@
Table S2: Refined crystal structure parameters for Cs3Fe2Cl9 at 1.6 K. The space group is $P6_{3}/mmc$ with lattice constants $a=7.1784(1)$ and $c=17.6369(2)$ Å. A uniform thermal parameter of $B_{\textrm{iso}}=0.1$ Å2 is assumed for all atoms. The goodness-of-fit parameters are $R_{p}=15.2\%$ and $R_{wp}=11.3\%$.
@@ V Inelastic neutron scattering experiments on CNCS
Inelastic neutron scattering (INS) experiments on powder samples of Cs3Fe2Cl9 were performed on CNCS Ehlers et al. (2011) at the SNS of the ORNL. About 5 g powder was sealed in an aluminum can in a helium filled glovebox. An orange cryostat was utilized to reach a base temperature of 5 K. Measurements were taken with incident neutron energies of $E_{i}=6.59$, 3.32, and 1.0 meV in the high flux chopper configuration at $T=2$ and 10 K. For each measuring condition, data were collected on an empty can and subtracted as background. Data reduction and projection were performed using the Mslice program in DAVE Azuah et al. (2009).
@@ VI Diffuse neutron scattering experiments on CORELLI
VII Neutron diffraction experiments on WAND2
@@
Single crystal neutron diffraction measurements were performed using the WAND2 diffractometer Frontzek et al. (2018) at the High Flux Isotope Reactor HFIR with $\lambda=1.486$ Å. A crystal with mass of $\sim 7.5$ mg was aligned with the $c$ axis vertical, and then sealed in an aluminum can with helium exchange gas. A vertical field cryomagnet was used, providing a base temperature of 1.5 K and a maximum field of $H=6$ T. Measurements were performed in phase I at $T$ = 1.5 K and no applied field rotating through 180° in 0.1° steps over 20 h. Measurements in phase II were made at $T$ = 1.5 K and $H=3.2$ T by rotating through 180° in 0.1° steps over 20 h. In phase III at $T=1.5$ K at $H=4.5$ T the sample was rotated through 90° in 0.1° steps over 10 h. In phase IV, at $T=1.5$ K and $H=5.5$ T the sample was rotated through 90° in 0.1° steps over 10 h. In phase V, at $T=4$ K and $H=5.5$ T the sample was rotated through 90° in 0.1° steps over 10 h. In phase VI, at $T=5.05$ K and $H=5.5$ T the sample was rotated over 90° in 0.1° steps with a total measurement time of 26 h.
@@
Figure S3: Single crystal neutron diffraction pattern for Cs3Fe2Cl9 in phases (a) III, (b) IV, and (c) V in the ($h$,kk,0) plane. The experimental conditions are listed in the text. The dashed rectangle emphasizes the area in which the relative intensity of the $\bm{q}_{\textrm{III}}$ and $2\bm{q}_{\textrm{III}}$ reflections are compared in the text.
@@
Figure S3 compares the neutron diffraction patterns in phases III, IV, and V. As explained in the main text, similar diffraction patterns were observed in these three phases. Especially, their magnetic propagation vector stays at the same $\bm{q}_{\textrm{III}}=\frac{5}{6}\times(\frac{1}{3},\frac{1}{3},0)$, with weak $2\bm{q}_{\textrm{III}}$ higher harmonics observed along the Brillouin zone boundary. Despite their similarities, careful comparisons of the diffraction patterns reveal some differences among these three phases. As emphasized by the dashed rectangle in each panel, the intensity of the $2\bm{q}_{\textrm{III}}$ reflection in phase III is weaker than that in phase IV, which suggests higher field-induced magnetization in phase IV. As compared in Fig. 4 of the main text for phases IV and V, the $\bm{q}_{\textrm{III}}$ reflection becomes weaker in phase V while the $2\bm{q}_{\textrm{III}}$ reflection stays almost unchanged. This evolution is consistent with a transition from a spiral structure in phase IV to a spin density wave (SDW) structure in phase V. For a spiral order like that in phase IV, both the in-plane and out-of-plane components contribute to the intensity of the main $\bm{q}_{\textrm{III}}$ reflections, while only the out-of-plane component, under renormalization as described in the section of ‘Analytical expressions for the field-induced phases’, contributes to the $2\bm{q}_{\textrm{III}}$ reflections. For the collinear SDW order in phase V, the in-plane component becomes zero, thus the intensity of the main $\bm{q}_{\textrm{III}}$ reflections will be strongly reduced compared that in phase IV.
VIII Refinements of the nuclear and magnetic structures
@@
Figure S5 summarizes the fitting results of the magnetic reflections collected on WAND2 in phases I (a), III (b), IV (c), and V (d). The measuring conditions are listed in the main text. In phase VI, only two nonequivalent magnetic reflections were obtained for each magnetic domain, which does not allow a reliable analysis of the magnetic structure. For phase I, the collinear magnetic structure as shown in Fig. 1(a) of the main text was assumed in the fits. The fitted magnitude of the ordered magnetic moment is 2.9(2) $\mu\rm{{}_{B}}$, with goodness-of-fit parameters $R_{F2}=20.3\%$ and $R_{F2w}=24.4\%$. For phases III-V, a helical magnetic structure was assumed based on our classical Monte Carlo simulations. The magnitudes of the in-plane magnetic moment, $M_{\parallel}$, and out-of-plane magnetic moment, $M_{\perp}$, were treated as fitting parameters. For phase III, the fitted magnitudes are $M_{\parallel}=1.9\pm 1.5$ and $M_{\perp}=1.9\pm 0.4$ $\mu\rm{{}_{B}}$, with goodness-of-fit parameters $R_{F2}=13.0\%$ and $R_{F2w}=14.7\%$. For phase IV, the fitted magnitudes are $M_{\parallel}=2.1\pm 1.6$ and $M_{\perp}=1.8\pm 0.5$ $\mu\rm{{}_{B}}$, with $R_{F2}=9.0\%$ and $R_{F2w}=10.1\%$. For phase V, the fitted magnitudes are $M_{\parallel}=1.8\pm 4.1$ and $M_{\perp}=1.2\pm 0.8$ $\mu\rm{{}_{B}}$, with $R_{F2}=3.2\%$ and $R_{F2w}=3.7\%$. Considering the standard deviations, the fitted magnitude of the in-plane moment $M_{\parallel}$ is consistent with our classical Monte Carlo simulations, where phases III and IV exhibit an elongated helical structure, and phase V exhibits a spin density wave structure with only the $S_{z}$ components.
IX Temperature evolution of the diffraction pattern
Figure S6 presents the temperature evolution of the neutron diffraction pattern for Cs3Fe2Cl9 in the ($h$,kk,0) plane. Experiments were measured on the WAND2 ($T=1.5$ K) and CORELLI ($T=6$ and 10 K) diffractometers using the same piece of crystal. The enhanced diffuse scattering intensity at $T=6$ K confirms its magnetic origin, which is replaced by magnetic Bragg peaks at $T=1.5$ K.
Figure S6: Temperature evolution of the neutron diffraction pattern in the ($h$,kk,0) plane. The pattern in panel (a) was measured on the WAND2 diffractometer. The patterns in (b) and (c) were measured on the CORELLI diffractometer, and data collected at $T=20$ K has been subtracted as background.
X Modeling the spin dynamics
To model the magnon dispersion observed in the CNCS experiments, we performed linear spin wave theory (LSWT) calculations using the SpinW software package Toth and Lake (2015). The powder-averaged dynamical structure factor, $S(Q,\omega)$, was calculated in steps of $0.05$ Å-1 over a momentum transfer range of $Q=0.35$ to $2.05$ Å-1 for the spectra collected with $E_{i}=3.32$ meV, and over the range of $Q=0.4$ to $0.95$ Å-1 for the spectra collected with $E_{i}=1.0$ meV. The calculated spectrum was convoluted with the theoretical instrumental energy resolution for each $E_{i}$ configuration plus an additional Gaussian broadening of 0.05 meV that accounts for thermal broadening. To reduce the impacts of neutron absorption, data below and above the energy transfer of $\Delta E=1.25$ meV in the $E_{i}=3.32$ meV spectra were treated separately with a fitted intensity ratio of about 3:2. The optimal exchange parameters were determined through a simultaneous fit for all three datasets including the low- and high-energy sections from the $E_{i}=3.32$ meV spectra and the $E_{i}=1$ meV spectra, using a weighting factor of 1:3:5, respectively.
Figure S7: Fit to the INS data using the $J_{1\textrm{-}5}$-$D_{z}$ model with (a) $E_{\mathrm{i}}=$ 3 meV data and (b) $E_{\mathrm{i}}=$ 1 meV data.
We began the INS fit with an initial $J_{123}$-$D_{z}$ model, which is the minimum model to realize spiral spin-liquids on the lattice of Cs3Fe2Cl9:
$\mathcal{H}=\sum_{n=1,2,3}J_{n}\bm{S}_{i}\cdot\bm{S}_{j}+D_{z}(S_{z})^{2}\textrm{.}$ (S1)
Starting from this minimum model that includes exchange interactions up to the third neighbors, we progressively includes further-neighbor exchange interactions to check whether they can further lower the $\chi^{2}$ factor. Through this procedure, we found that interactions up to the fourth neighbors, i.e., $J_{1}$ through $J_{4}$, are essential for accurately reproducing the INS spectra. The fifth-neighbor interactions, $J_{5}$, were also included in our model, as it slightly reduces the $\chi^{2}$ from 20.66 to 20.29, although its fitted strength of 0.0013(1) meV is already much weaker than the other terms. Further perturbations up to the seventh neighbors yield no improvement for the fit, and their fitted strengths, when included, are all in the orders of $10^{-4}$ meV or even less. We also investigated the impacts of symmetry-allowed anisotropic terms, including the Dzyaloshinskii-Moriya (DM) interactions or off-diagonal exchange terms, but no improvement was observed after introducing these terms in the fitting program. Therefore, in the final $J_{1\textrm{-}5}$-$D_{z}$ model, we considered only the isotropic exchange interactions up to the fifth neighbors. The fitting results for the INS spectra at representative $Q$ positions are summarized in Fig. S7.
XI Classical Monte Carlo simulations
@@
Classical Monte Carlo simulations for the $J_{1\textrm{-}5}$-$D_{z}$ model were performed using the SpinMC code that implements the single spin flip Metropolis algorithm bue . Unless otherwise specified, a $16\times 16\times 4$ supercell with 4096 spins was employed in simulations. Observables, including the heat capacity and magnetization, were averaged over $10^{5}$ of measurement sweeps after $2\times 10^{5}$ thermalization sweeps, where each sweep represents 8192 attempted spin flips at randomly selected sites. An overrelaxation sweep was applied after each Monte Carlo sweep to reduce the autocorrelation. Magnetic structure factors were calculated through fast Fourier transform using the FFTW package fft . The parallel tempering algorithm was utilized to facilitate thermal equilibrium, which was performed simultaneously over 80 replicas on 80 cores with a geometric series of temperatures between 1 and 10 K. After every 10 Monte Carlo sweeps, a replica exchange was attempted, and the successful exchange rates are above $\sim 25\%$ for all the neighboring replicas.
XII Phase diagram for the $J_{1\textrm{-}5}$-$D_{z}$ model